. XX company assembles electrical components for the last 10 days the company had averaged 9 rejected with standard deviation of 2 rejects.DD company averaged 8.5 rejected with standard deviation of 1.5 rejects. At the 0.05 level of significance can we conclude that there is more variation in the number of rejects per day for the XX company. 



1
Expert's answer
2021-12-27T16:21:16-0500

"H_0:\sigma_1=\sigma_2" , there is no more variation in the number of rejects per day for the XX company

"H_a:\sigma_1>\sigma_2" , there is more variation in the number of rejects per day for the XX company


test statistic:

"F=\frac{s_1^2}{s_2^2}=\frac{2^2}{1.5^2}=1.7778"

"df=n-1=10-1=9"


critical value:

"F_{(9,9)}=3.1789"


Since "F<F_{(9,9)}" we accept null hypothesis. There is no more variation in the number of rejects per day for the XX company.